# Proportional Representation in Metric Spaces and Low-Distortion Committee Selection

### 16 Dec 2023 v1

@@ 4.2 Stronger Proportional Fairness: The Key Lemma
[…]
@@ 6 Related Work
Representing a large point set is also a — sometimes explicit, sometimes implicit — goal of clustering points in a metric space. A review of this very large literature is beyond the scope of our work. Of the common objective functions (most notably, $k$-center, $k$-means, and $k$-median), $k$-median (Arya et al., 2001) is closest to our objective here, since it minimizes the sum of distances of points to the respective closest selected $k$ cluster centers. In the basic definition, like proportionally fair clustering, it suffers from the fact that a large and dense cluster can be “served” by just one selected representative.
Capacitated versions address this issue by limiting the number of points “served” by a cluster center (see, e.g., Byrka et al. (2015)). Our algorithms for both ordinal and cardinal models similarly assign to each representative $c$ a subset of size $p$ from $V$ that $c$ “covers”. To appreciate the difference, suppose that roughly $p$ points are at a large distance $M$ from all others points which are within a unit circle. If a representative is at distance $(1-\epsilon)M$ from these remote points, a $k$-median solution might include it since the objective reduction by $\epsilon M$ would outweigh any decisions made within the unit circle. In contrast, our proportional representation objective might exclude it since the relative improvement would be marginal. Thus, the proportional representation objective is much less sensitive to few outliers. It should also be noted that both objectives require resource augmentation for positive results (Byrka et al., 2015). Nonetheless, it would be interesting to explore the link between these two objectives and determining if one implies non-trivial guarantees for the other.
@@
Several papers have proposed extensions of the notion of (metric) distortion to multi-winner (or committee) elections. Goel et al. (2018) consider the cost of a set $R$ to be the sum of all distances between $V$ and the members of $R$. Because this sum decomposes, it is minimized by choosing the $k$ candidates individually closest to $V$ (in terms of sums of distances); as a result, the set $R$ tends to be as “homogeneous” as possible. For example, if $V$ is partitioned into two clusters of 51% and 49%, and there are enough candidates near each cluster, then the optimal solution would choose all candidates close to the cluster comprising 51% of $V$. We note that Goel et al. (2018) did not propose this objective as a natural objective for evaluating how representative $R$ is; rather, the goal was to show that many results relating (and bounding) distortion and fairness naturally generalize from single-winner to multi-winner elections. An alternative notion was proposed by Caragiannis et al. (2022). Their definition is parametrized by $q\leq k$: each individual $v\in V$ evaluates the cost of $R$ as the cost of the ${q}^{\rm th}$ closest representative in $R$; the objective to minimize is then the sum of these costs. When $q=1$, the objective coincides with (uncapacitated) $k$-median; in contrast, for $q=k$, a committee has low cost for $v$ only if all of its members are close to $v$. Thus, in the regime of large $q$, the definition suffers from the same drawback as that of Goel et al. (2018): it rewards committees $R$ (almost) all of whose members are close to the largest cluster within $V$. Caragiannis et al. (2022) show that while the objective can be well approximated with ordinal information in the “homogeneous” case $q>k/2$, the distortion is unbounded for $q<k/3$; though no results are shown under resource augmentation.
@@
There has been prior work relating notions of fairness, multi-winner elections, distortion, and the core. For example, Ebadian et al. (2022) study the utilitarian distortion of randomized single-winner voting rules. They use fairness both as a tool to derive novel low-distortion randomized voting rules, and — in a different notion of proportional fairness — as an optimization goal in its own right. They show that $\alpha$-approximate proportional fairness in their definition implies membership in the $\alpha$-core.
@@
In general cooperative games, the notion of core captures a stability desideratum: that the outcome be stable against deviations by any subgroup of players seeking better utilities. The exact coalitions and available “outside options” of utilities give rise to specific core solution concepts. This concept is further extended to inverse utility games, where the objective is to reduce the cost incurred by the players Tamir (1993); Chen et al. (2019). The stability found in the core of cost-sharing games closely mirrors the notion of proportional representation of points in a metric space. In a vein similar to proportional representation, the core in cost-sharing games guarantees that large fractions of players have low cost under the chosen solution, deterring them from deviating to alternatives.
In social choice, core definitions often emphasize proportionality, suggesting that a $\theta$ fraction of the population should “control” an equal fraction of the outcome (Moulin, 2003). In the context of electing a committee, various notions of core stability require that every sufficiently cohesive and large voter group should sufficiently “approve” of the committee. One branch of the literature explores such notions in the context of approval ballots, i.e., each voter either approves or disapproves each individual candidate. Depending on the precise definition of “cohesiveness” and “approval”, many versions of proportional representation emerge (Aziz et al., 2017; Sánchez-Fernández et al., 2017; Skowron, 2021) (see also the survey by Lackner and Skowron (2023)), which can be seen as stability or fairness measures. With ranked ballots, these definitions can be adapted by having voters approve their top $\tau$ choices, with an adjustable parameter $\tau$. This notion lies at the heart of expanding approvals. Another research direction, illustrated by the work of Cheng et al. (2020); Jiang et al. (2020), moves beyond expanding approvals, by defining the core for ranked ballots through a lexicographical ordering or costs associated with a (hidden) metric space. Chen et al. (2019); Li et al. (2021) utilize these core definitions to explore fair clustering, which are key focal points of this paper and are extensively discussed throughout.
@@ A.2 Lower Bound for small $\alpha$
[…]
@@ A.3 Separation between Cardinal and Ordinal Proportionally Fair Clustering
[…]

### 24 Jan 2024 v2

@@ 4.2 Stronger Proportional Fairness: The Key Lemma
We show that the committee $R$ returned by Algorithm 1 satisfies a somewhat stronger notion of proportional fairness, i.e., a modification of Definition 2.6. Lemma 4.7 shows that for any coalition $S$, the representatives $R[S]$ are already sufficiently attractive that $S$ will not unanimously deviate. The guarantee differs from Definition 2.6 only in the slight strengthening of replacing $R$ with $R[S]$.
@@ 6 Related Work
6.1 Multi-Winner Metric Distortion
@@
Several papers have proposed extensions of the notion of (metric) distortion to multi-winner (or committee) elections. Goel et al. (2018) consider the cost of a set $R$ to be the sum of all distances between $V$ and the members of $R$. Because this sum decomposes, it is minimized by choosing the $k$ candidates individually closest to $V$ (in terms of sums of distances); as a result, the set $R$ tends to be as “homogeneous” as possible. For example, if $V$ is partitioned into two clusters of 51% and 49%, and there are enough candidates near each cluster, then the optimal solution would choose all candidates close to the cluster comprising 51% of $V$. We note that Goel et al. (2018) did not propose this objective as a natural objective for evaluating how representative $R$ is; rather, the goal was to show that many results relating (and bounding) distortion and fairness naturally generalize from single-winner to multi-winner elections.
An alternative notion was proposed by Caragiannis et al. (2022). Their definition is parametrized by $q\leq k$: each individual $v\in V$ evaluates the cost of $R$ as the cost of the ${q}^{\rm th}$ closest representative in $R$; the objective to minimize is then the sum of these costs. When $q=1$, the objective coincides with (uncapacitated) $k$-median; in contrast, for $q=k$, a committee has low cost for $v$ only if all of its members are close to $v$. Thus, in the regime of large $q$, the definition suffers from the same drawback as that of Goel et al. (2018): it rewards committees $R$ (almost) all of whose members are close to the largest cluster within $V$. Caragiannis et al. (2022) show that while the objective can be well approximated with ordinal information in the “homogeneous” case $q>k/2$, the distortion is unbounded for $q<k/3$; though no results are shown under resource augmentation.
This notion of $q$-cost was used as a definition of fairness or representativeness in the context of sortition in subsequent work by Ebadian et al. (2022b) and Ebadian and Micha (2023). In Ebadian et al. (2022b) and part of Ebadian and Micha (2023), this notion is used primarily as an analysis tool: the analysis focused on uniformly random selection of individuals, and the representativeness/fairness of the resulting panels, i.e., no knowledge of the metric whatsoever (not even ordinal) was assumed. This is because the random selection procedure itself was a primary focus of the work. However, in addition, Ebadian and Micha (2023) were also interested in the question of how much more representativeness of panels could be achieved with full knowledge of the underlying metric; we discuss those results and their relationship with our work in Section 6.3.
6.2 The Core and Proportionality in Social Choice
@@
There has been a large amount of prior work relating notions of fairness, multi-winner elections, distortion, and the core. For example, Ebadian et al. (2022a) study the utilitarian distortion of randomized single-winner voting rules. They use fairness both as a tool to derive novel low-distortion randomized voting rules, and — under a different notion of proportional fairness — as an optimization goal in its own right. They show that $\alpha$-approximate proportional fairness in their definition implies membership in the $\alpha$-core.
@@
In general cooperative games, the notion of core captures a stability desideratum: that the outcome be stable against deviations by any subgroup of players seeking better utilities. The exact coalitions and available “outside options” of utilities give rise to specific core solution concepts. This concept is further extended to inverse utility games, where the objective is to reduce the cost incurred by the players (Tamir, 1993; Chen et al., 2019). The stability found in the core of cost-sharing games closely mirrors the notion of proportional representation of points in a metric space. In a vein similar to proportional representation, the core in cost-sharing games guarantees that large fractions of players have low cost under the chosen solution, deterring them from deviating to alternatives.
In social choice, core definitions often emphasize proportionality, suggesting that a $\theta$ fraction of the population should “control” an equal fraction of the outcome (Moulin, 2003). In the context of electing a committee, various notions of core stability require that every sufficiently cohesive and large voter group should sufficiently “approve” of the committee. Early research by Dummett (1984) aimed to ensure proportionality for solid coalitions (PSC) on ranked ballots. These coalitions consist of voter groups whose set of top candidates are the same (for corresponding set sizes), though their ordering of these top candidates may differ.
Subsequently, more research focused on approval ballots instead of ranked ballots; here, each voter specifies for each candidate whether they approve of the candidate or not; equivalently, the voter specifies the set of all approved candidates. Aziz et al. (2017) introduced the concept of Justified Representation (JR). Justified Representation (JR) requires that cohesive groups be well-represented in the committee as follows: if a set of $\theta n$ voters is cohesive in the sense of having at least $\theta k$ candidates whom they all approve, then at least $\theta k$ of those unanimously approved candidates must be selected. Unfortunately, a committee achieving JR may not exist. Subsequent work of Sánchez-Fernández et al. (2017); Peters et al. (2021); Skowron (2021) further refined this concept, in part with an eye towards ensuring existence of a committee satisfying the definition. They consider various alternate and subtly differing definitions of cohesiveness; see also the survey by Lackner and Skowron (2023).
In studying axiomatic approaches for proportional representation, it has been observed repeatedly (see the in-depth discussion of this strand of literature in Brill and Peters (2023)) that various cohesiveness definitions for coalitions — both for approval and ranked ballots — are very demanding. As a result, only few coalitions are cohesive, imposing few requirements on a selected committee. In response, Brill and Peters (2023) propose new definitions for both approval and ranked ballots. Their PJR+ notion requires the following. For any coalition $S$ which, based on proportionality, is “entitled” to $\ell$ committee members, if the union of committee members approved by at least one voter in the coalition has size smaller than $\ell$, then the committee is not allowed to omit any candidate who is unanimously approved by the coalition. The rank-PJR+ notion for ranked ballots requires this to hold for the “approval ballots” obtained by truncating all ranked ballots at position $r$, for all $r$. This constitutes a more demanding notion of proportional representation under ranked choice ballots, compared to PSC (Dummett, 1984). Brill and Peters (2023) show that while Single-Transferable Vote fails rank-PJR+, the Expanding Approvals Rule satisfies it.
General notions of stability are also studied by Cheng et al. (2020); Jiang et al. (2020). Cheng et al. (2020) again consider deviation threats in which coalitions $S$ may deviate to a set of size $\frac{|S|}{n}\cdot k$ if there exists such a set that they prefer over the chosen committee. They consider both approval and ranked ballots: for approval ballots, voters prefer the committee that contains more candidates whom the voter approves, while for ranked ballots, voters prefer the committee with the single highest-ranked candidate according to the voter’s order. The main focus of Cheng et al. (2020) is on lotteries over committees that achieve stability in this sense (since no deterministic stable committee may exist). Jiang et al. (2020) consider more general monotone preferences of voters over committees. They consider approximate stability, similar to our resource augmentation: they require that no coalition $S$ have a deviation threat of size $\frac{|S|}{cn}\cdot k$ for $c\geq 1$ that they prefer over the chosen committee. The main contribution of Jiang et al. (2020) is to show that for a constant $c$, there always exists an approximately stable committee in this sense, so long as the voters’ preferences are monotone. They prove this by starting from a lottery over committees, which is then derandomized.
6.3 Clustering, Fairness and Proportionality
Representing a large point set is also a — sometimes explicit, sometimes implicit — goal of clustering points in a metric space. A review of this very large literature is beyond the scope of our work. Of the common objective functions (most notably, $k$-center, $k$-means, and $k$-median), $k$-median (Arya et al., 2001) is closest to our objective here, since it minimizes the sum of distances of points to the respective closest of the selected $k$ cluster centers. In the basic definition, like proportionally fair clustering, it suffers from the fact that a large and dense cluster can be “served” by just one selected representative.
Capacitated versions address this issue by limiting the number of points “served” by a cluster center (see, e.g., Byrka et al. (2015)). Our algorithms for both ordinal and cardinal models similarly assign to each representative $c$ a subset of size $p$ from $V$ that $c$ “covers”. To appreciate the difference between the objectives, suppose that roughly $p$ points are at a large distance $M$ from all others points which are within a unit circle. If a representative is at distance $(1-\epsilon)M$ from these remote points, a $k$-median solution might include it since the objective reduction by $\epsilon M$ would outweigh any decisions made within the unit circle. In contrast, our proportional representation objective might exclude it since the relative improvement would be marginal. Thus, the proportional representation objective is much less sensitive to few outliers. It should also be noted that both objectives require resource augmentation for positive results (Byrka et al., 2015). While the objectives look quite different, it would be interesting to explore a link between them and to determine if one implies non-trivial guarantees for the other.
Notions of fairness have also been incorporated into clustering problems more explicitly. These can be roughly divided into fairness notions towards individuals and towards coalitions. A notion of fairness towards individuals was proposed by Jung et al. (2020); Chakrabarti et al. (2022): they consider a committee of size $k$ to be $\alpha$-fair if for each individual $i$, the committee contains a member that is no more than $\alpha$ times as far from $i$ compared to $i$’s ${n/k}^{\rm th}$ closest candidate. Fairness in this sense may come at a significant cost for the overall utility function (such as $k$-median), and Mahabadi and Vakilian (2020); Vakilian and Yalçıner (2022) give algorithms with improved tradeoffs between the objectives of fairness and overall clustering quality.
An alternative notion of individual fairness was considered by Chakrabarti et al. (2022). Their notion somewhat resembles envy-freeness: each point has a self-declared set of “peers”, and would like its distance to its closest center to be larger by at most a factor $\alpha$ than that of its peers; the latter could be the minimum or average distance of peers to their closest center.
Yet another notion was considered in the work of Ebadian and Micha (2023). Their work was motivated by Sortition: choosing random citizen panels such that each citizen has (approximately) uniform probability of being part of the panel, and the panel overall is representative in the sense of multi-winner elections discussed above: i.e., for each individual, the panel contains $q$ members approximately closest to the individual. In order to obtain a good tradeoff between individual inclusion probabilities and approximation to the core, Ebadian and Micha (2023) consider another modified version of the Greedy Capture algorithm. Like our Truncated Greedy Capture algorithm, their version freezes a ball when representatives are chosen; however, this happens only when a ball contains $q$ times the Hare quota many points, and at this point, the $q$ representatives from the ball are chosen uniformly at random. Interestingly, the approximation to the core is shown to be the same $\frac{5+\sqrt{41}}{2}$ which appears in our analysis of the Expanding Approvals Rule for Proportionally Fair Clustering. The technical steps leading to this bound (i.e., specific triangle inequalities and combinations) are similar, even though the algorithms and overall analysis are different.
An alternative approach, more in line with our definition, is to require fairness (or lack of deviation incentives) for sufficiently large groups. Perhaps the most prominent examples are the notions of proportional fairness (Chen et al., 2019; Micha and Shah, 2020) and approximate core (Li et al., 2021), both discussed in depth in Section 2.2.
Simultaneously and independently of our work, Aziz et al. (2023) also propose a notion of proportional representation. Their motivation is quite similar to ours: they also observe that neither proportionally fair clustering nor the approximate core achieve the goal (stated in those works) of representing clusters proportionally to their size. They formulate a basic axiom called Unanimous Proportionality (UP), which requires that if there is a coalition $S$ comprising a $\theta$ fraction of the population all of whose members are in the exact same location, then a $\theta$ fraction of the committee should be made up of points closest to $S$. This is a clean axiomatization of the issue we discuss in Section 2.2; indeed, our notion of proportional representation also leads to satisfying the UP axiom. Aziz et al. (2023) then propose a definition they term Proportionally Representative Fairness (PRF). This notion requires the following, for each set $S$. Let $y$ be the diameter of $S$, and $\ell$ the number of candidates who are at distance at most $y$ from all members of $S$. Let $\theta$ the fraction of the population that $S$ comprises. Then, the committee must contain at least $\mathop{min}(\theta k,\ell)$ candidates who are at distance at most $y$ from all members of $S$. Aziz et al. (2023) then propose an algorithm for selecting a committee satisfying their definition. The algorithm is essentially an Expanding Approvals Rule with known distances; that is, balls around candidates are grown, and when a ball contains $n/k$ voters, the candidate is included, and the voters are “removed” (by having their weight decreased). As with the algorithms we consider, proportionality is ensured by having the balls around included candidates not grow any further. Notice that the algorithms are designed with knowledge of the metric space. The notions of proportional representation of our work and Aziz et al. (2023) are obviously similar in their goals and approaches, although we are not aware of a formal reduction. It should be noted that while our notion is a strict strengthening of the notion of core fairness proposed by (Li et al., 2021), this is not the case for the notion of Aziz et al. (2023).
In further simultaneous and independent work, Kellerhals and Peters (2023) aim to draw connections between the plethora of different fairness/proportionality notions alluded to above. Among others, they show that proportional fairness and individual fairness imply each other (up to losses of constant factors). More importantly, they define metric notions of ranked fairness. Recall from Section 6.2 that notions of justified representation (JR, EJR, PJR, PJR+) can be extended to ranked (rather than approval) ballots by considering all the cutoffs $r$, and requiring the conditions to hold for each implied approval ballot. As an alternative, Kellerhals and Peters (2023) propose notions of rank-JR, rank-PJR, rank-PJR+ based on distance cutoffs: they consider agents as approving all candidates within distance $r$, and require that the corresponding justified representation axioms hold for the implied approval ballots. They show that various natural algorithms (such as Greedy Capture and a distance-based Expanding Approvals Rule) satisfy corresponding axioms. Furthermore, they show implications between these new axioms and proportional and individual fairness (as well as the proportional representation notion of Aziz et al. (2023)), obtaining results on the representation guarantees for the outputs of these algorithms for various notions. In several cases, the obtained guarantees match the best known results.
@@ 7 Concluding Remarks and Open Questions
Acknowledgements
We would like to thank Evi Micha for useful discussions, and Jannik Peters for pointing us to additional related and simultaneous work. This work was supported under ARO MURI grant W911NF1810208.
@@ A.2 Lower Bound for small $\alpha$
$\frac{b_{i}}{p-b_{i}}$ $\geq\frac{\lfloor{n/(k+1)}\rfloor}{\left\lceil{n/k}\right\rceil-\lfloor{n/(k+1)}\rfloor}$
@@ A.3 Separation between Cardinal and Ordinal Proportionally Fair Clustering
In Corollary 4.3, we showed the existence of a deterministic 5.71-proportionally fair clustering algorithm under ordinal information. On the other hand, Chen et al. [2019], Micha and Shah [2020] gave a deterministic $1+\sqrt{2}\approx 2.41$-proportionally fair clustering algorithm when the metric space is known to the algorithm. Here, we show a separation between proportionally fair clustering in the ordinal and cardinal models, by giving an instance under ordinal information for which no deterministic algorithm can be better than $2+\sqrt{5}\approx 4.23$-proportionally fair.